### v1

@@ 7.1 Anderson-Chaplain (AC) model
The AC model employs an on-lattice biased random walk to simulate tip EC movement in response to local levels of VEGF and fibronectin. A regular grid of $N\times N$ points spaced at intervals of $h$ is placed on the square domain $\mathcal{I}$ (we fix $N=200$ and $h=0.05$), and tip ECs move through the domain one grid space at a time. The initial concentration of VEGF is as defined above and the initial concentration of fibronectin is $f(x,y,0)=1-y$. The time evolution of the VEGF and fibronectin concentrations, and the related probability that a tip EC moves left, right, down or up on the lattice, are derived from a system of Partial Differential Equations (PDEs).
@@
On the lattice, the VEGF and fibronectin initial conditions become $c^{0}_{l,m}=m/N$ and $f^{0}_{l,m}=1-m/N$. The discretisations (7)–(8) may then be used to update the VEGF and fibronectin concentrations $c^{t}_{l,m}$ and $f^{t}_{l,m}$ at grid locations $(l,m)$ in discrete time-steps of duration $\Delta t=0.01$ (using a discrete version of the no-flux boundary conditions). Rather than using (7) to compute the concentration of ECs at lattice points, the factors $P_{0},P_{1},P_{2},P_{3}$ and $P_{4}$, whose formulae are given in (9)–(13), are used to determine the probability that an individual EC makes a move on the square lattice.
@@
At each time step, each EC either remains at its location $(l,m)$, or moves to $(l-1,m)$, $(l+1,m)$, $(l,m-1)$ or $(l,m+1)$ according to the probabilities $\hat{P}_{0},\hat{P}_{1},\hat{P}_{2},\hat{P}_{3}$ and $\hat{P}_{4}$, where $\hat{P}_{j}=P_{j}/(P_{0}+P_{1}+P_{2}+P_{3}+P_{4})$ for $j=0,1,2,3,4$. To simulate this, a uniform random number $u\sim\mathcal{U}_{[0,1]}$ is drawn for each tip EC at each time-step. If $u\in[0,\hat{P}_{0})$ then $s^{t+\Delta t}=s^{t}$, if $u\in[\hat{P}_{0},\hat{P}_{1})$, the EC moves left, and so on. If this procedure prescribes a move outside of the domain $\mathcal{I}$, that tip EC terminates and is not considered for any further moves.
@@ 7.2 Stokes-Lauffenberger (SL) model
In the SL model (Stokes et al., 1991), tip ECs can move in any direction (off-lattice). The movement of each tip EC is governed by (14), a two-dimensional Stochastic Differential Equation (SDE) for EC velocity $\mathbf{V}(t)$. At discrete time-steps, Equation (14) is solved using the Euler-Maruyama method and Equation (15) is integrated via the forward Euler method to give EC position.
@@
In the SL model, the parameter $\beta$ models the cell’s resistance to movement, $\mathbf{W}(t)$ is a two-dimensional Wiener process (in which increments $\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$ are independently and normally distributed), and the parameter $\sigma$ represents an EC’s tendency to deviate from its current direction. $\mathbf{\Psi}(t)=\kappa\nabla c\sin\absolutevalue{\frac{\psi}{2}}$ | models movement due to chemotaxis, where $c$ is the VEGF concentration and $\psi$ is the angle that the current velocity $\mathbf{V}(t)$ makes with the direction of steepest increase in $c$. Given $c(x,y)=y$, the direction of increasing VEGF concentration is always $(0,1)^{T}$, which simplifies the chemotaxis term. The parameter $\kappa$ measures the strength with which the EC velocity re-orients up spatial gradients of $c$ (towards the tumour). ECs initialised along the bottom edge of the square domain are assigned a small initial velocity in the $y$-direction: $v^{0}=(0,0.5)^{T}$. The discretisation (16) of Equations (14) and (15) is then used to simulate EC velocity $v^{t}$ and position $s^{t}$ at each time-step.
@@ 7.3 Plank-Sleeman (PS) model
The PS model (Plank and Sleeman, 2004) assigns a constant speed $\hat{s}$ to each EC and varies the angle $\phi$ that a tip EC’s velocity vector makes with the horizontal direction. An EC’s position $s^{t}=(s^{t}_{x},s^{t}_{y})\in\mathcal{I}$ is then modelled by the system of ordinary differential equations:
$\frac{\text{d}s^{t}_{x}}{\text{d}t}=\hat{s}\cos{\phi},\ \ \ \ \frac{\text{d}s^{t}_{y}}{\text{d}t}=\hat{s}\sin{\phi}.$ ) . (17)
@@
$\mu(\phi)=-d_{c}\absolutevalue{\nabla c}\sin(\phi-\phi_{c}).$ (18)
In Equation (18), the turning coefficient $d_{c}$ determines how often an EC angle re-orients its movement angle towards the direction of increasing VEGF concentration $\phi_{c}=\pi/2$. Let $\hat{\tau}^{\pm}_{n}$ denote the probability that an EC rotates through a fixed angle of ±$\pm\hat{\phi}$ on the $n$-th time-step. It can be shown (Plank and Sleeman, 2004) that, if the mean turning rate is defined by (18), then $\hat{\tau}^{\pm}_{n}$ are given by:
@@ 7.5 Model simulation, schematic and parameters
We fix all but four parameter values in each angiogenesis model and simulate the movement of tip ECs for $400$ time-steps. See Table 1 for a list of the parameters that vary and their ranges. See also the schematic at the top of Figure 4 for an illustration of model movement rules.
We convert data simulated from the three angiogenesis models into a common format for data analysis. The common format should retain the spatial structure of simulated networks while discretising them to allow the computation of spatially-averaged and topological statistics. Considering a simulated network at its final timestep, we overlay a regular square grid of $K=200$ points spaced in intervals of $h=0.05$ in co-ordinate directions over the simulation domain. We then say an angiogenesis dataset is the point-cloud consisting of the $(x,y)$ locations of ECs within this discretised image. The AC model simulates data on such a grid already, and we convert data from the SL and PS models into the common format by populating those grid squares which intersect the (off-lattice) paths traced by tip ECs. This common format will also be applied to observed data and ensures that summary statistics give a fair comparison between data simulated by each model.
Figure 4: Schematic showing tip EC movement rules in the Anderson-Chaplain (AC), Stokes-Lauffenberger (SL) and Plank-Sleeman (PS) models. In the AC model, tip ECs move on a grid according to probabilities $\hat{P}_{j}=P_{j}/(P_{0}+P_{1}+P_{2}+P_{3}+P_{4})$ for $j=0,1,2,3,4$. Higher values of the chemotaxis and haptotaxis parameters induce a bias into those probabilities which specify a movement towards increasing concentrations of VEGF and fibronectin respectively. The SL model updates the velocity $v_{i}^{t}$ of the EC at location $s^{t}$ as a weighted sum of the current velocity $w$, randomness $r$ and chemotaxis $c$. Randomness and chemotaxis and parameters moderate the weight of the corresponding terms when updating the velocity. The PS model rotates the movement angle $\phi$ between an EC’s velocity vector and the horizontal by $\hat{\phi}$ with transition probabilities $\hat{\tau}_{n}^{+}$ and $\hat{\tau}_{n}^{-}$ and ECs move a fixed distance $\hat{s}$ in the new direction at each time-step. A turning rate parameter moderates how often the EC’s angle of movement updates, and a turning bias parameter changes how likely such an update is to favor the direction of increasing VEGF concentration. All models use the same rules for branching–a tip EC at location $s^{t}$ bifurcates into two tip ECs that move independently when $t$ is greater than the minimum age for branching parameter $a_{\text{br}}$ and the concentration of VEGF at $s^{t}$ is greater than the VEGF threshold for branching parameter $c_{\text{br}}$. If a movement rule would cause an EC to move into a location already occupied by an EC, that EC instead anastomoses and is considered for no further movement. Created in https://BioRender.com
@@
$[0,0.5]$ Taken from (Nardini et al., 2021)
@@
$[0,0.5]$ Taken from (Nardini et al., 2021)
@@ 8 Computation of Summary Statistics
8.1 Spatially-Averaged Statistics
An angiogenesis dataset $\mathcal{D}$ is a point-cloud consisting of $(x,y)$ grid locations which contain simulated EC at the final simulation timestep. We compute the mean, standard deviation, minimum, maximum, range, interquartile range, and the 10th, 25th, 75th and 90th percentiles of the $x$ and $y$-coordinates of points in $\mathcal{D}$. Concatenating these values gives a list of $20$ spatially-averaged statistics, $10$ in the horizontal ($x$) coordinate direction and $10$ in the vertical ($y$) coordinate direction.
@@ 8.2 Extended Persistent Homology
We detect and quantify topological structure in angiogenesis datasets using Extended Persistent Homology (EPH). Here, we give details of how we compute and vectorise EPH, giving a brief overview of how it arises from Persistent Homology (PH) and outlining our reasons for using EPH instead of PH.
To compute PH, spatial data is first converted to a nested sequence of simplicial complexes $\{\Sigma_{k}\}_{k=0}^{K}$ known as a filtration. A simplicial complex is a collection of vertices (00-simplices), where subsets of vertices can be connected by edges ($1$-simplices), triangles ($2$-simplices), tetrahedra ($3$-simplices) and their higher dimensional analogues. A face of a simplex is defined as a subset of its vertices along with the simplices that are connected to them. Simplicial complexes satisfy the property that a face of any simplex, or the intersection of two simplices, is also a simplex in the complex.
The dimension-$p$ PH of the filtration is then the sequence (21) of $\mathbb{F}$-vector spaces $H_{p}(\Sigma_{k})$ (we fix $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$) together with maps induced by the inclusion maps between the simplicial complexes. The basis elements of $H_{p}(\Sigma_{k})$ correspond to $p$-dimensional holes in $\Sigma_{k}$ (Carlsson, 2009), also referred to as topological features of the simplicial complex. Since each $\Sigma_{k}\subset\Sigma_{k+1}$, inclusion maps $\iota_{k}:\Sigma_{k}\to\Sigma_{k+1}$ map simplices in $\Sigma_{k}$ to their counterparts in the larger $\Sigma_{k+1}$. The inclusion maps $\iota_{k}$ then induce linear maps $\iota_{k}^{*}:H_{p}(\Sigma_{k})\to H_{p}(\Sigma_{k+1})$ in sequence (21) (by functoriality) which allow topological features to be tracked through the filtration.
@@
The Structure Theorem of PH (Zomorodian and Carlsson, 2005) states that sequence (21) uniquely decomposes into a direct sum of interval modules $I_{b,d}$, which are sequences of $\mathbb{F}$-vector spaces (22), where $\eta_{k}^{*}$ are identity maps when $b\leq k<d$ and zero maps otherwise.
$0\displaystyle 0\xrightarrow{\eta_{0}^{*}}0\dots 0\xrightarrow{\eta_{b-1}^{*}}\mathbb{F}\xrightarrow{\eta_{b}^{*}}\mathbb{F}\dots\mathbb{F}\xrightarrow{\eta_{d-1}^{*}}0\dots 0\xrightarrow{\eta_{K-1}^{*}}$ 0 (22)
Persistence pairs $(b,d)$ that define interval modules (22) correspond to topological features in the filtration. The birth $b$ is the first index $k$ for which the corresponding $p$-dimensional hole appears in the filtration (and the index $k$ in (22) for which the corresponding topological feature is not in the image of $\eta_{k-1}^{*}$). The death $d$ is the filtration index $k$ where the $p$-dimensional hole is filled in by additional simplices (and the index $k$ in (22) for which the corresponding topological feature is mapped to 00 by $\eta_{k-1}^{*}$). The persistence $d-b$ measures how long the corresponding topological feature persists in the filtration. Persistence pairs may be plotted as points in birth-death coordinates in a Persistence Diagram (PD). PDs are stable (Chazal et al., 2009) to small perturbations in the underlying point-cloud, making them useful topological summaries of spatial data.
Nardini et al. (2021) applied PH to study angiogenesis datasets $\mathcal{D}$ by constructing a sweeping-plane filtration as follows. First, $\mathcal{D}$ is converted to a simplicial complex $\Sigma$. Each point $(x,y)$ in the angiogenesis dataset is represented in $\Sigma$ by a vertex (00-simplex). If two vertices represent two grid locations in $\mathcal{D}$ which are adjacent (in the Moore neighbourhood), they are connected with edges ($1$-simplices). Collections of three edges are connected with triangle ($2$-simplex) if each pair of edges shares a vertex. A function $f:\Sigma\to\mathbb{R}$ is then defined such if a vertex $u$ represents a point $(x,y)\in\mathcal{D}$, $f(u)=y$. The value of $f$ on other simplices within $\Sigma$ is then simply the maximum of the values of $f$ on their vertices. The sweeping-plane filtration is then defined as the sublevel sets $\Sigma_{k}=f^{-1}(-\infty,k/K]$, where $k=0,1,\dots,K$ and $K$ is the resolution of the image (e.g. $K=200)$.
@@
Since $H_{p}(\Sigma_{K})=H_{p}(\Sigma_{K},\Sigma^{K})$, the sequences (21) and (23) may be concatenated to give a single sequence (of length $12k+$ 1), which is known as the dimension-$p$ EPH of the filtration. The structure theorem applies to this sequence, meaning that topological features may be extracted as persistence pairs $(b,d)$. Intuitively, the sweeping-plane filtration scans through $\Sigma$ from bottom to top and (21) detects topological features which are found below $k/K$. The complexes $\Sigma_{K}/\Sigma^{k}$ then scan back down from top to bottom, collapsing all simplices above $k/K$ to a single point, with the sequence (23) detecting topological features in the resulting complexes. In particular, all simplices merge in $\Sigma_{K}/\Sigma^{0}$, meaning that no persistence pairs extracted from (21)–(23) have $d=\infty$.
Persistence pairs $(b,d)$ representing topological features which are born and die within the sequence (21) are called ordinary persistence pairs. The sweeping-plane filtration detects such topological features in dimension $p=0$ when a path of ECs anastomoses with another (without forming a loop). Topological features which are born and die in the sequence (23) are called relative persistence pairs. In dimension $p=1$ such topological features result from branch points where an EC bifurcates in two separate paths which do not later anastomose to form loops. Those topological features which are born in (21) but die in (23) are extracted as extended persistence pairs. If $b<d$ and $p=0$, extended persistence pairs represent connected components–distinct EC paths which remain separated. If $b>d$ and $p=1$, extended persistence pairs represent loops in the simulated vasculature.
We use two functions $f$ to create two sweeping-plane filtrations from an angiogenesis dataset $\mathcal{D}$. Creating a simplicial complex $\Sigma$ as above, we define $f_{\text{horizontal}},f_{\text{vertical}}:\Sigma\to\mathbb{R}$ on vertices $u$ representing points $(x,y)\in\mathcal{D}$ by $f_{\text{horizontal}}(u)=x$, $f_{\text{vertical}}(u)=y$, and on other simplices as before. Persistence pairs resulting from the sequences (21)–(23) then represent the size and location of topological features measured in the horizontal ($x$) and vertical ($y$) coordinate directions. Persistence pairs may be plotted on an Extended Persistence Diagram (EPD), an example of which is given in Figure 5.
8.3 Topological Statistics

### v2

@@ 7.1 Anderson-Chaplain (AC) model
The AC model employs an on-lattice biased random walk to simulate tip EC movement in response to local levels of VEGF and fibronectin. A regular grid of $N\times N$ points spaced at intervals of $h$ is placed on the square domain $\mathcal{I}$ (we fix $N=200$ and $h=0.05$), and tip ECs move through the domain one grid space at a time. The initial concentration of VEGF is as defined above and the initial concentration of fibronectin is f(x,y,0)=1yf(x,y,0)=1-y. The time evolution of the VEGF and fibronectin concentrations, and the related probability that a tip EC moves left, right, down or up on the lattice, are derived from a system of Partial Differential Equations (PDEs).
@@
On the lattice, the VEGF and fibronectin initial conditions become $c^{0}_{l,m}=m/N$ and $f^{0}_{l,m}=1-m/N$. The discretisations (5)–(6) may then be used to update the VEGF and fibronectin concentrations $c^{t}_{l,m}$ and $f^{t}_{l,m}$ at grid locations $(l,m)$ in discrete time-steps of duration $\Delta t$ (using a discrete version of the no-flux boundary conditions). Rather than using (5) to compute the concentration of ECs at lattice points, the factors $P_{0},P_{1},P_{2},P_{3}$ and $P_{4}$, whose formulae are given in (7)–(11), are used to determine the probability that an individual EC makes a move on the square lattice.
@@
At each time step, each EC either remains at its location $(l,m)$, or moves to (l1,m)(l-1,m), $(l+1,m)$, (l,m1)(l,m-1) or $(l,m+1)$ according to the probabilities $\hat{P}_{0},\hat{P}_{1},\hat{P}_{2},\hat{P}_{3}$ and $\hat{P}_{4}$, where $\hat{P}_{j}=P_{j}/(P_{0}+P_{1}+P_{2}+P_{3}+P_{4})$ for $j=0,1,2,3,4$. To simulate this, a uniform random number $u\sim\mathcal{U}_{[0,1]}$ is drawn for each tip EC at each time-step. If $u\in[0,\hat{P}_{0})$ then $s^{t+\Delta t}=s^{t}$, if $u\in[\hat{P}_{0},\hat{P}_{1})$, the EC moves left, and so on. If this procedure prescribes a move outside of the domain $\mathcal{I}$, that tip EC terminates and is not considered for any further moves.
@@ 7.2 Stokes-Lauffenberger (SL) model
In the SL model (stokesLauffenberger), tip ECs can move in any direction (off-lattice). The movement of each tip EC is governed by (12), a two-dimensional Stochastic Differential Equation (SDE) for EC velocity $\mathbf{V}(t)$. At discrete time-steps, Equation (12) is solved using the Euler-Maruyama method and Equation (13) is integrated via the forward Euler method to give EC position.
@@
In the SL model, the parameter $\beta$ models the cell’s resistance to movement, $\mathbf{W}(t)$ is a two-dimensional Wiener process (in which increments $\mathbf{W}(t+\Delta t)-\mathbf{W}(t)$ are independently and normally distributed), and the parameter $\sigma$ represents an EC’s tendency to deviate from its current direction. $|\mathbf{\Psi}(t)=\kappa\nabla c\sin\absolutevalue{\frac{\psi}{2}}$ models movement due to chemotaxis, where $c$ is the VEGF concentration and $\psi$ is the angle that the current velocity $\mathbf{V}(t)$ makes with the direction of steepest increase in $c$. Given $c(x,y)=y$, the direction of increasing VEGF concentration is always $(0,1)^{T}$, which simplifies the chemotaxis term. The parameter $\kappa$ measures the strength with which the EC velocity re-orients up spatial gradients of $c$ (towards the tumour). ECs initialised along the bottom edge of the square domain are assigned a small initial velocity in the $y$-direction: $v^{0}=(0,0.5)^{T}$. The discretisation (14) of Equations (12) and (13) is then used to simulate EC velocity $v^{t}$ and position $s^{t}$ at each time-step.
@@ 7.3 Plank-Sleeman (PS) model
The PS model (plank_sleeman) assigns a constant speed $\hat{s}$ to each EC and varies the angle $\phi$ that a tip EC’s velocity vector makes with the horizontal direction. An EC’s position $s^{t}=(s^{t}_{x},s^{t}_{y})\in\mathcal{I}$ is then modelled by the system of ordinary differential equations:
[…]
@@
[…]
In Equation (16), the turning coefficient $d_{c}$ determines how often an EC angle re-orients its movement angle towards the direction of increasing VEGF concentration $\phi_{c}=\pi/2$. Let $\hat{\tau}^{\pm}_{n}$ denote the probability that an EC rotates through a fixed angle of $\pm\hat{\phi}$ on the $n$-th time-step. It can be shown (plank_sleeman) that, if the mean turning rate is defined by (16), then $\hat{\tau}^{\pm}_{n}$ are given by:
@@ 7.5 Model simulation, schematic and parameters
In each model, we use a timestep of $\Delta t=0.01$, and simulate EC movement over the (dimensionless) time interval $\mathcal{T}=[0,t_{\text{final}}]$. We wish to fix $t_{\text{final}}$ to allow a fair comparison between models, and large enough to ensure we observe enough EC movement to infer model parameters. ECs move approximately 0.005 spatial units during each timestep in each model–this is the exactly the grid spacing in the AC model, and we scaled the SL and PS models so that ECs move approximately this distance in each timestep too. If all ECs moved directly upwards in each timestep, they would reach the tumour at time $t=2$. We choose $t_{\textnormal{final}}=4$, thus simulating each model for $400$ timesteps, which will allow most meandering or bifurcated ECs to terminate by either reaching the tumour, anastomosing, or moving outside of the simulation domain.
We fix all but four parameter values in each angiogenesis model–see Table 1 for a list of the parameters that vary and their ranges. See also the schematic at the top of Figure 5 for an illustration of model movement rules.
Figure 5: Schematic showing tip EC movement rules in the Anderson-Chaplain (AC), Stokes-Lauffenberger (SL) and Plank-Sleeman (PS) models. In the AC model, tip ECs move on a grid according to probabilities $\hat{P}_{j}=P_{j}/(P_{0}+P_{1}+P_{2}+P_{3}+P_{4})$ for $j=0,1,2,3,4$. Higher values of the chemotaxis and haptotaxis parameters induce a bias into those probabilities which specify a movement towards increasing concentrations of VEGF and fibronectin respectively. The SL model updates the velocity $v_{i}^{t}$ of the EC at location $s^{t}$ as a weighted sum of the current velocity $w$, randomness $r$ and chemotaxis $c$. Randomness and chemotaxis and parameters regulate the weight of the corresponding terms when updating the velocity. The PS model rotates the movement angle $\phi$ between an EC’s velocity vector and the horizontal by $\hat{\phi}$ with transition probabilities $\hat{\tau}_{n}^{+}$ and $\hat{\tau}_{n}^{-}$ and ECs move a fixed distance $\hat{s}$ in the new direction at each time-step. A turning rate parameter regulates how often the EC’s angle of movement updates, and a turning bias parameter changes how likely such an update is to favor the direction of increasing VEGF concentration. All models use the same rules for branching–a tip EC at location $s^{t}$ bifurcates into two tip ECs that move independently when $t$ is greater than the minimum age for branching parameter $a_{\text{br}}$ and the concentration of VEGF at $s^{t}$ is greater than the VEGF threshold for branching parameter $c_{\text{br}}$. If a movement rule would cause an EC to move into a location already occupied by an EC, that EC instead anastomoses and is considered for no further movement. Created in https://BioRender.com
@@
$[0,0.5]$ Taken from (nardiniclustering)
@@
$[0,0.5]$ Taken from (nardiniclustering)
@@ 8 Computation of Summary Statistics
We convert data simulated from the three angiogenesis models (Anderson-Chaplain (AC), Stokes-Lauffenberger (SL), Plank-Sleeman (PS)) into a common format and then compute summary statistics.
Although simulations describe the movement of ECs in the domain $\mathcal{I}$ throughout the time interval $0,t_{\text{final}}]$, EC trails do not remodel after they have been laid down in the models we study. The value $t_{\text{final}}=4$ we choose allows the majority of ECs to reach the tumour within the simulation time (see the previous section for details). Therefore, the snapshot of the domain at the final timestep contains most of the information provided by the full simulation. Hence, and in order to simplify calculations, we consider only this final timestep snapshot when computing summary statistics.
The common format retains the spatial structure of simulated networks while discretising them to allow the computation of spatially-averaged and topological summary statistics. Considering a simulated network at its final timestep, we overlay a regular square grid of $K=200$ points spaced in intervals of $h=0.05$ in co-ordinate directions. We then say an angiogenesis dataset $\mathcal{D}$ is the point-cloud consisting of the $(x,y)$ locations of ECs within this discretised image. The AC model simulates data on such a grid already, and we convert data from the SL and PS models into the common format by populating those grid squares which intersect the (off-lattice) paths traced by tip ECs. This common format will also be applied to observed data and it ensures that summary statistics give a fair comparison between data simulated by each model.
8.1 Spatially-Averaged Summary Statistics
An angiogenesis dataset $\mathcal{D}$ is a point-cloud consisting of $(x,y)$ grid locations which contain simulated EC at the final simulation timestep. We compute the mean, standard deviation, minimum, maximum, range, interquartile range, and the 10th, 25th, 75th and 90th percentiles of the $x$ and $y$-coordinates of points in $\mathcal{D}$. Concatenating these values gives a list of $20$ spatially-averaged summary statistics, $10$ in the horizontal ($x$) coordinate direction and $10$ in the vertical ($y$) coordinate direction.
@@ 8.2 Extended Persistent Homology
We detect and quantify topological structure in angiogenesis datasets using extended persistent homology (EPH). Here, we give details of how we compute and vectorise EPH, outlining how it arises from persistent homology (PH) and explaining our reasons for using EPH instead of PH.
To compute PH, spatial data is first converted to a nested sequence of simplicial complexes $\{\Sigma_{k}\}_{k=0}^{K}$ known as a filtration. A simplicial complex is a collection of vertices (0-simplices), where subsets of vertices can be connected by edges ($1$-simplices), triangles ($2$-simplices), tetrahedra ($3$-simplices) and their higher dimensional analogues. A face of a simplex is defined as a subset of its vertices along with the simplices that are connected to them. Simplicial complexes satisfy the property that a face of any simplex, or the intersection of two simplices, is also a simplex in the complex.
The dimension-$p$ PH of the filtration is then the sequence (19) of $\mathbb{F}$-vector spaces $H_{p}(\Sigma_{k})$ (we fix $\mathbb{F}=\mathbb{Z}/2\mathbb{Z}$) together with maps induced by the inclusion maps between the simplicial complexes. The basis elements of $H_{p}(\Sigma_{k})$ correspond to $p$-dimensional holes in $\Sigma_{k}$ (carlssonTDA), also referred to as topological features of the simplicial complex. Since each $\Sigma_{k}\subset\Sigma_{k+1}$, inclusion maps $\iota_{k}:\Sigma_{k}\to\Sigma_{k+1}$ map simplices in $\Sigma_{k}$ to their counterparts in the larger $\Sigma_{k+1}$. The inclusion maps $\iota_{k}$ then induce linear maps $\iota_{k}^{*}:H_{p}(\Sigma_{k})\to H_{p}(\Sigma_{k+1})$ in sequence (19) (by functoriality) which allow topological features to be tracked through the filtration.
@@
The Structure Theorem of PH (structurePH) states that sequence (19) uniquely decomposes into a direct sum of interval modules $I_{b,d}$, which are sequences of $\mathbb{F}$-vector spaces (20), where $\eta_{k}^{*}$ are identity maps when $b\leq k<d$ and zero maps otherwise.
$0\xrightarrow{\eta_{0}^{*}}0\dots 0\xrightarrow{\eta_{b-1}^{*}}\mathbb{F}\xrightarrow{\eta_{b}^{*}}\mathbb{F}\dots\mathbb{F}\xrightarrow{\eta_{d-1}^{*}}0\dots 0\xrightarrow{\eta_{K-1}^{*}}0$ (20)
Persistence pairs $(b,d)$ that define interval modules (20) correspond to topological features in the filtration. The birth $b$ is the first index $k$ for which the corresponding $p$-dimensional hole appears in the filtration (and the first index $k$ in (20) for which the corresponding topological feature is in the image of $\eta_{k}^{*}$ but not in the image of $\eta_{k-1}^{*}$). The death $d$ is the filtration index $k$ where the $p$-dimensional hole is filled in by additional simplices (and the first index $k$ in (20) for which the corresponding topological feature is mapped to 0 by $\eta_{k-1}^{*}$). The persistence dbd-b measures how long the corresponding topological feature persists in the filtration. Persistence pairs may be plotted as points in birth-death coordinates in a persistence diagram (PD). PDs are stable (PHstable) to small perturbations in the underlying point-cloud, making them useful topological summaries of spatial data.
nardiniclustering applied PH to study angiogenesis datasets $\mathcal{D}$ by constructing a sweeping-plane filtration as follows. First, $\mathcal{D}$ is converted to a simplicial complex $\Sigma$. Each point $(x,y)$ in the angiogenesis dataset is represented in $\Sigma$ by a vertex (0-simplex). If two vertices represent two grid locations in $\mathcal{D}$ which are adjacent (in the Moore neighbourhood), they are connected with edges ($1$-simplices). Collections of three edges are connected with triangle ($2$-simplex) if each pair of edges shares a vertex. A function $f:\Sigma\to\mathbb{R}$ is then defined such if a vertex $u$ represents a point $(x,y)\in\mathcal{D}$, $f(u)=y$. The value of $f$ on other simplices within $\Sigma$ is then simply the maximum of the values of $f$ on their vertices. The sweeping-plane filtration is then defined as the sublevel sets $\Sigma_{k}=f^{-1}(-\infty,k/K]$, where $k=0,1,\dots,K$ and $K$ is the resolution of the image (e.g. $K=200)$.
@@
Since $H_{p}(\Sigma_{K})=H_{p}(\Sigma_{K},\Sigma^{K})$, the sequences (19) and (21) may be concatenated to give a single sequence (of length $2k+1$), which is known as the dimension-$p$ EPH of the filtration. The structure theorem applies to this sequence, meaning that topological features may be extracted as persistence pairs $(b,d)$. Intuitively, the sweeping-plane filtration scans through $\Sigma$ from bottom to top and the vector spaces (19) detect topological features which are found below $k/K$. The complexes $\Sigma_{K}/\Sigma^{k}$ then scan back down from top to bottom, collapsing all simplices above $k/K$ to a single point, with the sequence (21) detecting topological features in the resulting complexes. In particular, all simplices merge in $\Sigma_{K}/\Sigma^{0}$, and no persistence pairs extracted from the combined sequence (19)–(21) have $d=\infty$.
All features computed through EPH are born and die at some point in the concatenated sequence (19)–(21). Features may be classified into four types depending depending on which part of this sequence they are born and die in. If a feature is born and dies within sequence (19), it is called ordinary. If a feature is born and dies within in sequence (21), it is called relative. If a feature is born in sequence (19) and dies in sequence (21) it is called extended. Such persistence pairs are called extended+ if $b>d$ and extended- if $b<d$.
We use two functions $f$ to create two sweeping-plane filtrations from an angiogenesis dataset $\mathcal{D}$. After converting $\mathcal{D}$ into a simplicial complex $\Sigma$ as above, we define $f_{\text{horizontal}},f_{\text{vertical}}:\Sigma\to\mathbb{R}$ on vertices $u$ representing points $(x,y)\in\mathcal{D}$ by $f_{\text{horizontal}}(u)=x$, $f_{\text{vertical}}(u)=y$, and on other simplices as before. Using these filtrations to compute EPH, persistence pairs resulting from the sequences (19)–(21) then represent the size and location of topological features measured in the horizontal ($x$) and vertical ($y$) coordinate directions. Persistence pairs may be plotted on an extended persistence diagram (EPD). Figure 6 contains an example computation of EPDs using the horizontal and vertical sweeping-plane filtration. Each EPD contains one feature of each type (ordinary, relative, extended+, and extended-). This Figure is an expanded version of Figure 2 of the main text, and it also shows the extra information gained by computing EPH over PH.
8.3 Topological Summary Statistics